OpenStudy (anonymous):

What is the margin of error for a yes/no survey of 200 people with a proportion of 65% who answered "yes"? A. ±3.37% B. ±5.06% C. ±6.75%D. ±8.43% E. ±10.12% *please help*

OpenStudy (amistre64):

sqrt(pq/n)

OpenStudy (amistre64):

\[\sqrt{\frac{.65(1-.65)}{200}}\] \[\sqrt{\frac{.65(.35)}{200}}\]

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

youre welcome, its one of the formulas that i do tend to remember :)